OpenStudy (anonymous):

An orchestra consists of 25 members. The youngest member is 32 years old. She leaves and is replaced by a new member who is 30 years old.whaat is effect on average age , mean??

OpenStudy (anonymous):

\[\huge\frac{ 32+x_2+....+x25 }{ 25 }=y\]

OpenStudy (anonymous):

previously the average age is (x+32)/25..where x is the sum of ages of other 24 members..now i think you can tell what is the average age after 32 year member is replaced by 30 year old member..

OpenStudy (anonymous):

\[\huge \frac{ 30+x_2+.....+x25 }{ 25 }=w\]

OpenStudy (anonymous):

\[\huge \frac{25w+2}{ 25 }=y\]

OpenStudy (anonymous):

\[\huge w=\frac{ 25y-2 }{ 25 }\]

OpenStudy (anonymous):

the mean of their ages will decrease by 0.08 ??

OpenStudy (anonymous):

median??

Directrix (directrix):

If A is the average age of the 25 people before the 32-year old left, then 25*A is the sum of the ages. After the 32-year old left and a 30-year old joined, then (25*A - 2)/25 is the new average age. (25*A - 2)/25 = 25*A/25 - 2/25 = A - 2/25 = A - 8/100 = A - .08. So, does the average go down by .08 people? If so, I agree with @deena123

OpenStudy (kropot72):

Let x = original average age. Original sum of ages = 25x New sum of ages = 25x - 2 New mean is \[\frac{25x-2}{25}=x-0.08\]

Directrix (directrix):

> median?? @non008 Are you asking a second question?

OpenStudy (anonymous):

Tell me about median ?

OpenStudy (anonymous):

no median of above question.

Directrix (directrix):

I did the average problem on a smaller case with 3 people and the same problem constraints as posted. The mean went down by 2/3 a point. So, I'm conjecturing that under the same conditions but with 100 people in the group, the mean would drop by 2/100 a unit. Just saying.

OpenStudy (kropot72):

The median will not change when the age of the youngest player is reduced. The median is the middle value in a ranked set of data values.

OpenStudy (anonymous):

thanks @Directrix and @kropot72

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

Weights of male mountain lions follow the normal distribution with a median of 140 lb and an interquartile range of 8 lb. How to find the 55th percentile??

OpenStudy (anonymous):

2A : 2.0 POINTS Find the 75th percentile of the weights.

OpenStudy (anonymous):

PROBLEM 3 : 2.0 POINTS Undergraduates in a large university class were asked to record the time they spent on the most recent homework. In a histogram of the recorded times, the height of the bar over the interval “1 hour to 3 hours” is 4% per hour. The height of the bar over the interval “3 hours to 6 hours” is 2% per hour. If the histogram is redrawn with just a single bar over the interval “1 hour to 6 hours”, what should the height of the bar be? [Calculate your answer in % per hour, but do not enter the units of measurement.]

OpenStudy (anonymous):

The scatter diagram of midterm scores and final scores in a large class is football shaped. For about 80% of the students, the regression estimate of final score based on midterm score is correct to within 15 points. For about 95% of the students, the regression estimate of final score based on midterm score is correct to within ___________ points.